OpenStudy (anonymous):

Assuming e^x equals y, and ln y =x, then would it be appropriate to say Dx(ln y) = Dx(x)?

hartnn (hartnn):

if you meant \[\dfrac{d}{dx}\ln y=\dfrac{d}{dx}x\] then yes.